Compute the first thrre derivatives of \\(f(x)=2x^{5}+x^{\\frac{3}{2}}-\\frac{1}{2x}\\)
1
Expert's answer
2019-09-19T12:26:04-0400
 Answer

"f'''(x)=120x^2-{\frac{3}{8}}x^{-\frac{3}{2}}+3x^{-4}."

Explanation

According the linearity property of the derivative for any triple of constants c1, c2, c3 and three functions f, g and h we have

(c1f + c2g + c3h)’ = c1f’ + c2g’ + c3h’.

On the other hand the differentiation rule of power function can be written in the form

"(x^{n})'=nx^{n-1}"

Because,"\frac{1}{x}=x^{-1}" we can write

"f(x)=2x^5+x^{\frac{3}{2}}-\frac{1}{2}x^{-1}"

So,

"f'(x)=(2x^5+x^{\frac{3}{2}}-\frac{1}{2}x^{-1})'=10x^4+{\frac{3}{2}}x^{\frac{1}{2}}+\frac{1}{2}x^{-2}"

For the second derivative we get

"f''(x)=(10x^4+{\frac{3}{2}}x^{\frac{1}{2}}+\frac{1}{2}x^{-2})'=40x^3+{\frac{3}{4}}x^{-\frac{1}{2}}-x^{-3}"

And for the third derivative we have

"f'''(x)=(40x^3+{\frac{3}{4}}x^{-\frac{1}{2}}-x^{-3})'=120x^2-{\frac{3}{8}}x^{-\frac{3}{2}}+3x^{-4}."